OpenStudy (anonymous):

Find P(A or B) if the events are disjointed. P(A) = 6/25, and P(B) = 9/25

OpenStudy (anonymous):

p(a)+p(b)

OpenStudy (anonymous):

if the events are are disjoint the its is P(A or B) =P(A) + P(B) = 6/25 +9/25 =15/25=3/5=0.6